mixture of isomeric Diphenyl and Dichloro Tetrachloroethane is used as a thermic
fluid in a liquid phase heating system. The thermic fluid enters an indirect fired heater
at a temperature of 450 K and leaves the heater at a temperature of 550 K. Calculate the
supply of heat to the heater per kg of the liquid heated. The heat capacity of the fluid is
given by the equation
CP = 1.435 + 2.19 × 10–3T, kJ/(kg K)
1
Expert's answer
2020-11-12T06:30:24-0500

"CP_{450} = 1.435 + 2.19 \times 10^{-3} \times 450 = 2.4205 \;kJ/kg\;K"

"CP_{550} = 1.435 + 2.19 \times 10^{-3} \times 550 = 2.6395 \;kJ/kg\;K"

"∆CP = 2.6395 – 2.4205 = 0.219 \;kJ/kg\;K"

The heat capacity was changed by 0.219 units, so we need to supply 0.219 kJ of heat to the heater per kg of the liquid heated.